# From 2D Images to 3D Model: Weakly Supervised Multi-View Face Reconstruction with Deep Fusion

### v4

@@ 4.1 Photo Loss
4.2 Mask Loss
Photo loss focuses on the pixel difference between two pictures. It is difficult to constrain the size of the facial feature area in the two pictures. For example, the nose color is very similar to that of the cheeks, thereby leading to difficulties for the photo loss to notice the boundary line between them. For this reason, we introduce mask loss to narrow the facial features of the input image and the rendered image. The division and labeling of the facial features are shown in Fig. 3. According to BiSeNet (Yu et al., 2018) we adopt cross entropy to calculate the mask loss $L_{m}$:
$L_{m}=$ $-\frac{1}{N_{p}\mathcal{V}}\sum_{v=1}^{V}\sum_{p=1}^{N_{p}}(y_{p}\log\left(\hat{y_{p}}\right)$ (3)
$+\left(1-y_{p}\right)\log\left(1-\hat{y_{p}}\right)),$
where $N_{p}$ is number of pixels in face mask. $\hat{y_{p}}$ donates the category prediction of pixels in face mask and $y_{p}$ is the label.
4.3 Landmark Loss
We also adopt 2D landmarks and 3D landmarks for weakly supervised training. We use 3D face alignment method (Bulat and Tzimiropoulos, 2017) to generate 68 landmarks $\left\{l_{n}\right\}$ as the ground-truth. Then the corresponding points in the predicted 3D face point cloud are projected to 2D as predicted 2D landmarks $\left\{l_{n}^{\prime}\right\}$. Then the multi-view 2D landmark loss can be calculated:
$L_{l\_2d}=\frac{1}{NV}\sum_{v=1}^{V}\sum_{n=1}^{N}\omega_{n}\left\|l_{n}^{v}-l_{n}^{v^{\prime}}\right\|_{2},$ (4)
where $\omega_{n}$ is the weight for each landmark. We set the weight to 20 only for the nose and inner mouth landmarks, and to 1 otherwise.
2D landmarks are still insufficient for the reconstruction of 3D face shapes. In order to obtain better reconstruction effect, we select 101 3D landmarks $\left\{q_{n}^{\prime}\right\}$ to impose a weak constraint on the shape of the 3D face. According to the 3DMM index, 101 predicted landmarks $\left\{q_{n}\right\}$ can be found. Then, we select 7 points $\left\{a_{n}^{\prime}\right\}$ and $\left\{a_{n}\right\}$ in $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$ respectively as alignment points to calculate the alignment parameters of $\left\{q_{n}^{\prime}\right\}$ and $\left\{q_{n}\right\}$. The alignment parameters include: scale $\mathbf{s}$, rotation $\mathbf{R}$ and translation $\mathbf{t}$. These parameters can be obtained by the following optimization equation (Sanyal et al., 2019):
$Optim(\mathbf{s},\mathbf{R},\mathbf{t}=\min_{\mathbf{s},\mathbf{R},\mathbf{t}}\sum_{i}\left\|\mathbf{a}_{i}^{\prime}-\mathbf{s}\left(\mathbf{R}\cdot\mathbf{a}_{i}+\mathbf{t}\right)\right\|_{2}.$ (5)
After the optimal $\mathbf{s}$, $\mathbf{R}$ and $\mathbf{t}$ are obtained, the predicted 101 landmarks $\left\{q_{n}\right\}$ can be converted to the space of $\left\{q_{n}^{\prime}\right\}$ as $\left\{q_{nt}\right\}=s\left(\mathbf{R}\cdot\mathbf{q}_{n}+\mathbf{t}\right)$.
Then the multi-view 3D landmark loss can be calculated:
$L_{l\_3d}=\frac{1}{N}\sum_{n=1}^{N}\left\|q_{nt}-q_{n}^{{\prime}}\right\|_{2}.$ (6)
In summary, the landmark loss can be expressed as:
$L_{l}=\omega_{2d}L_{l\_2d}+\omega_{3d}L_{l\_3d},$ (7)
where $\omega_{2d}$ and $\omega_{3d}$ represent respectively the weight of 2D landmark loss and 3D landmark loss. In this work, we set them to $0.02$ and $1$ as tuned empirically.
4.4 Regularization Loss
To suppress the generation of distorted faces, we add the regularization loss which is commonly-used in face reconstruction task (Thies et al., 2016; Tewari et al., 2018; Deng et al., 2019; Shang et al., 2020):
$L_{\text{reg}}=\omega_{\alpha}\|\boldsymbol{\alpha}\|^{2}+\omega_{\beta}\|\boldsymbol{\beta}\|^{2}+\omega_{\gamma}\|\boldsymbol{\gamma}\|^{2},$ (8)
where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ are 3DMM parameter vectors that the network predicted. $\omega_{\alpha}$, $\omega_{\beta}$ and $\omega_{\gamma}$ are the weights for 3DMM parameter vectors. Following Deep3DFace (Deng et al., 2019), we set them to $1$, $0.8$ and $0.017$ with fine tuning.
4.5 Overall Loss
@@
$\begin{split}L_{all}=\omega_{p}L_{p}+\omega_{m}L_{m}+\omega_{l}L_{l}+\omega_{reg}L_{reg},\end{split}$
where $\omega_{p}$, $\omega_{m}$, $\omega_{l}$, $\omega_{reg}$ are the weights for photo loss, mask loss, landmark loss and regularization loss. Following Deep3DFace, we set $\omega_{reg}=3.0\times 10^{-4}$. Since $\omega_{2d}$ and $\omega_{3d}$ have been determined, we adjust other parameters as $\omega_{l}=1$, $\omega_{p}=4$ and $\omega_{m}=3$ respectively.
@@
Note that 3D landmarks is not a must for our method. For general weakly supervised methods, we do not incorporate $L_{l\_3d}$ for comparison (Deng et al., 2019; Shang et al., 2020). Since semi-supervised methods HRN (Lei et al., 2023) adopts the entire 3D scans to attain training label, we add $L_{l\_3d}$ for comparative analysis.
5 Experiment
5.1 Implementation Details
@@
Dataset. Following BiseNet (Yu et al., 2018), we adopt CelebAMask-HQ (Lee et al., 2020) to pretrain Face Parse Net. Furthermore, Pixel-Face (Lyu et al., 2020) and Multi-PIE (Gross et al., 2008) are introduced to provide 3D landmarks label and multi-view faces. Pixel-Face contains $855$ subjects ranging in age from $18$ to $80$ years old. Each subject has $7$ or $23$ samples of different expressions. Pixel-Face has 3D mesh file of each sample as ground-truth but not 3DMM parameters or angle of multi-view images. In the experiment, the training/test split was set to $0.8$. For Multi-PIE, we leverage the version provided by TP-GAN (Huang et al., 2017), which will reduce a lot of data preprocessing work. This dataset contains about 140k multi-view images and does not provide 3D mesh and landmarks. We utilize the method of STNet (Chen et al., 2016) to detect and align faces. Then each input image is cropped and resized to 224$\times$224.
Network. Our network structure is shown in Fig. 2 and described in the methodology section. Based on the pre-trained BiseNet (Yu et al., 2018) with frozen weights, the face parse network is located in the beginning and end of the network. We implement DF-MVR by pytorch (Paszke et al., 2019). It is trained on 2 NVIDIA RTX 3090 GPUs card with a batch size of 16 for 512 epochs. We adopt Adam as the gradient descent method. The initial learning rate is set to 0.0001 which decays with the cosine anneal schedule. The average inference time of DF-MVR is 244ms.
Evaluation Metric. Following the previous works, RMSE (mm) (Wu et al., 2019; Deng et al., 2019; Shang et al., 2020) is used to compute point-to-plane L2 distance between predict 3D scans and ground-truth 3D scans. Concretely, the front face area is cropped for metrics calculation instead of using a complete BFM model (Sanyal et al., 2019; Deng et al., 2019; Shang et al., 2020). Before calculating point-to-plane L2 distance, the predicted 3D scans need to be registered with ground-truth 3D scans. Also, we take the ICP registration method (Li et al., 2017), the same as Deep3DFace (Deng et al., 2019).
5.2 Comparison to SOTAs
@@
We compare our method with the existing weakly supervised MVRs on both Pixel-Face and Bosphorus datasets. More specifically, only the four methods with codes can be found in the literature related to multi-view weakly supervised 3D face reconstruction. MGCNet (Shang et al., 2020) takes multiple images for training, and then uses one single image for testing. We select the best results among the three images for display. Deep3DFace (Deng et al., 2019) does not release their source codes of its scoring network. Therefore, we use their codes to train/test on Pixel-Face to select the best results among the three images. Moreover, MVFR (Xiao et al., 2022) necessitates scale_RT data, rendering it currently effective only for the Facescape dataset (Yang et al., 2020; Zhu et al., 2023). HRN (Lei et al., 2023) does not release its training code, but it provides a pretrained model based on numerous data. To this end, we use its pretrained model and source code for multi-view face reconstruction. Given the scarcity of open sources for this task, we release our full codes (training & test) and pretrained model in GitHub [^4] https://github.com/weiguangzhao/DF_MVR.
Table 1: Comparison to SOTAs. $\dagger$ stands for the necessity to leverage 3D ground-truth (GT) data for semi-supervised training. ${}^{\dagger}$ utilizes the entire 3D ground-truth scan to attain both deformation and displacement maps. ${}^{\dagger}$ only takes a sparse set of 3D landmarks.
@@
Deep3DFace (Deng et al., 2019) Pixel-Face 1.6641
MGCNet (Shang et al., 2020) Pixel-Face 1.8877
@@
Deep3DFace (Deng et al., 2019) Bosphorus 1.4777
MGCNet (Shang et al., 2020) Bosphorus 1.4418
@@
${}^{\dagger}$ (Lei et al., 2023) Pixel-Face 1.4061
${}^{\dagger}$ Pixel-Face 1.4040
Comparsion on Pixel-Face. The quantitative results of the comparison are shown in Table 1. Since Deep3DFace (Deng et al., 2019) and MGCNet (Shang et al., 2020) did not use 3D landmarks, to be fair, we also provide the results of our model without using 3D landmarks for comparison. As observed, Our model (without 3D landmarks) shows $5.2\%$ improvement compared to the existing methods. While HRN (Lei et al., 2023) adopts the complete 3D ground-truth scan to attain deformation map and displacement map, we only take very light annotation, 101 3D landmarks to constraint the face reconstruction. ${}^{\dagger}$ shows very competitive results with fewer 3D annotation, compared to ${}^{\dagger}$.
The visual comparison is shown in Fig. 5 given 3-view faces. It is evident that our predicted model is more accurate, especially in terms of facial depth estimation in the facial features.
@@
Comparsion on Bosphorus. Since Pixel-Face dataset is collected based on Asian faces, the samples it contains are limited by regions. On the other hand, Pixel-Face dataset is only made under one lighting environment and three fixed camera angles. Taking into account the diversity, we adopt the richer Multi-PIE dataset for training and Bosphorus dataset for testing. Both datasets contain faces across continents and are collected in various lighting environments and shooting angles. Because Multi-PIE does not provide 3D landmarks, we remove $L_{l\_3d}$ to train DF-MVR for this part. The tradition delaunay algorithm is utilized to generate mesh files from point clouds for Bosphorus dataset. We set the left and right inputs to 45 degrees for comparison. More angle robustness experiments can be later seen in Table 4. The quantitative results of the comparison are shown in Table 1. Our model (without 3D landmarks) exhibits $3.0\%$ improvement compared to the existing methods. The qualitative comparison is also provided in Fig. 6.
@@
5.3 Single-view Reconstruction.
The single-view reconstruction method only requires one image to generate the 3D face. From the practical viewpoint, it is more flexible though it may be inferior to multi-view methods in terms of accuracy. Our method can also be adapted in the single view scenario. More specifically, during the training process, we only change the input, without changing other parts. As shown in Fig. 7, the original input has been changed to four different forms, according to the probability: $\boldsymbol{P}_{\mathrm{a}}^{\prime}$, $\boldsymbol{P}_{\mathrm{l}}$, $\boldsymbol{P}_{\mathrm{f}}$,$\boldsymbol{P}_{\mathrm{r}}$. The input of multi-view images still needs to be dominant to preserve accuracy, so we set its probability to 2/3, and the other inputs equally distribute with the probability of 1/3.
Figure 7: Inputs for Single-view Reconstruction. ’None’ means a black image.
The parameterized results of the comparison are shown in Table 2. As observed, our proposed model also attains the superior performance. In the case of single image testing, our model is more effective than Deep3DFace and MGCNet.
Table 2: Comparison of RMSE (mm)
Method Dataset RMSE (mm) $\downarrow$
Deep3DFace (Deng et al., 2019) Pixel-Face 1.6641
MGCNet (Shang et al., 2020) Pixel-Face 1.8877
Ours (single-view) Pixel-Face 1.5437
5.4 Ablation Study
In order to verify the effectiveness of MulEn-Unet and the mask mechanism we designed, we perform more ablation experiments as shown in Table 3. The mean and standard deviation of RMSE are again used as the evaluation metric. First, from v1, v2, v6, it can be found that the multi-view feature fusion network we designed is superior to traditional CNN and Unet in this task. Then, the results of v2 and v6 hint that the multi-layer feature interaction in the feature extraction stage is better than the direct concatenation of features at the end. To be fair, we set the number of layers of RedNet and ResNet to 50. Through the RMSE of v3 and v6, it is clear that RedNet performs better than ResNet in this task. For v4, we not only remove the mask loss but also the face mask $\mathbf{I}_{A}$, $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$, which are concatenated to the original image. By comparing v5 and v6, we can see that the face mask mechanism promotes the network to generate a higher-precision model. Moreover, we verify the face mask and mask loss separately. Mask loss has a much smaller effect on the accuracy of the model than face mask. While the qualitative result of mask loss is positive for the adjustment of the model during training, we retain this part for reference. Finally, we remove $L_{lan\_3d}$, which means that our model can be trained with only three multi-view images without any 3D label (as denoted as v5). The result also shows that our model is accurate and stable.
Table 3: Ablation Study
@@ 5.4 Ablation Study
v1 CNN RedNet Y Y 1.5415
v2 Unet RedNet Y Y 1.5207
v3 MulEn-Unet ResNet Y Y 1.5206
v4 MulEn-Unet RedNet N Y 1.4656
v5 MulEn-Unet RedNet Y N 1.5770
v6 MulEn-Unet RedNet Y Y 1.4040
5.5 Robustness Analysis
Multi-PIE and Bosphorus are collected by different cameras. DF-MVR uses Multi-PIE for training and achieves good test results on Bosphorus, which implies that our model has a certain robustness to camera device changes. Moreover, we further explore robustness of DF-MVR for shooting angles. We fix the $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$ face angles while adjusting the $\mathbf{I}_{A}$ face angle. Table 4 and Fig. 8(a) show that our model consistently achieve the best results in different shooting angles, implying that our model exhibits the best robustness. Random means we choose the degree from $10^{\circ}$, $20^{\circ}$, $30^{\circ}$ and $45^{\circ}$ randomly. The light environment is also random in this experiment.
Table 4: Robustness Analysis on Bosphorus dataset.
@@ 5.5 Robustness Analysis
Deep3DFace (Deng et al., 2019) 1.5030 1.4975 1.4924 1.4777 1.4905
MGCNet  (Shang et al., 2020) 1.4419 1.4412 1.4413 1.4418 1.4416
@@
We also validate the angle robustness of our model visually. These results are plotted in Fig. 8(a). Detailed error maps are introduced to display the error for each facial area directly. Take the second sample of Fig. 8(b) for a typical example where we fix $\mathbf{I}_{B}$ and $\mathbf{I}_{c}$ while changing the face angle of $\mathbf{I}_{A}$ from $10^{\circ}$ to $45^{\circ}$. Obviously, the error maps generated from these angles do not change much. Furthermore, we provide the error maps of Deep3DFace (Deng et al., 2019) for comparison. At different angles, DF-MVR reconstructs the face with less error, especially in critical areas (e.g. eye, brow, nose and mouth).
(a) Quantitative results on Bosphorus dataset.
(b) Reconstruction Error Maps. The errors are expressed from small to large in the following order: blue, green, yellow, red.
Figure 8: Robustness Analysis.
6 Conclusion
In this paper, we design a novel end-to-end weakly supervised multi-view 3D face reconstruction network that exploits multi-view encoding to a single decoding framework with skip connections, able to extract, integrate, and compensate deep features with attention. In addition, we develop a multi-view face parse network to learn, identify, and emphasize the critical common face area. Combining pixelwise photometric loss, mask loss, and landmark loss, we complete the weakly supervised training. Extensive experiments verify the effectiveness of our model.
7 Ethics Statement
In this paper, we presented visualizations of the results for Pixel-Face and Bosphorus dataset. Specifically, we have only showcased Pixel-Face facial information for a single individual, and this information has been obtained with the explicit consent of the individual and the Pixel Face dataset team, allowing for publication. Additionally, we presented Bosphorus facial information for three individuals, all of whom are included in the publicly disclosed roster of the Bosphorus dataset. Our approach leverages multi-view images to generate high-fidelity 3D facial models, meeting the demands of industries such as film, gaming, virtual characters, and more. Due to the strong privacy implications associated with facial data, we suggest establishing detailed rules to regulate the use of facial reconstruction.

### v5

@@ IV-A Photo Loss
IV-B Mask Loss
Photo loss focuses on the pixel difference between two pictures. It is difficult to constrain the size of the facial feature area in the two pictures. For example, the nose color is very similar to that of the cheeks, thereby leading to difficulties for the photo loss to notice the boundary line between them. For this reason, we introduce mask loss to narrow the facial features of the input image and the rendered image. The division and labeling of the facial features are shown in Fig. 2. According to BiSeNet [15] we adopt cross entropy to calculate the mask loss $L_{m}$:
$L_{m}=-\frac{1}{N_{p}\mathcal{V}}\sum_{v=1}^{V}\sum_{p=1}^{N_{p}}(y_{p}\log\left(\hat{y_{p}}\right)+\left(1-y_{p}\right)\log\left(1-\hat{y_{p}}\right)),$ (5)
where $N_{p}$ is the number of pixels in a face mask. $\hat{y_{p}}$ donates the category prediction of pixels in face mask and $y_{p}$ is the label.
IV-C Overall Loss
@@
$L_{all}=\omega_{p}L_{p}+\omega_{m}L_{m}+\omega_{l}L_{l}+\omega_{reg}L_{reg},$ (6)
[…]
@@
Note that 3D landmarks are not a must for our method. For general weakly supervised methods, we do not incorporate $L_{l\_3d}$ for comparison [10, 11]. Since semi-supervised methods HRN [24] adopt the entire 3D scans to attain training labels, we add $L_{l\_3d}$ for comparative analysis.
V Experiment
V-A Experiment settings
@@
Dataset and Evaluation Metric. Following BiseNet [15], we adopt CelebAMask-HQ [35] to pretrain Face Parse Net. Furthermore, Pixel-Face [16] and Multi-PIE [36] are introduced to provide 3D landmarks labels and multi-view faces. Pixel-Face contains $855$ subjects ranging in age from $18$ to $80$ years old. Each subject has $7$ or $23$ samples of different expressions. Pixel-Face has the 3D mesh file of each sample as ground truth but not 3DMM parameters or angle of multi-view images. Additionally, following the previous works, RMSE (mm) [10, 11] is used to compute point-to-plane L2 distance between predicted 3D scans and ground-truth 3D scans.
V-B Comparison to SOTAs
@@
We compare our method with the existing weakly supervised MVRs on both Pixel-Face and Bosphorus datasets. More specifically, only the four methods with codes can be found in the literature related to multi-view weakly supervised 3D face reconstruction. MGCNet [11] takes multiple images for training and then uses one single image for testing. We select the best results among the three images for display. Deep3DFace [10] does not release the source codes of its scoring network. Therefore, we use their codes to train/test on Pixel-Face to select the best results among the three images. HRN [24] does not release its training code, but it provides a pre-trained model based on numerous data. To this end, we use its pre-trained model and source code for multi-view face reconstruction. The comparison results have been provided in Tab. I. $\dagger$ stands for the necessity to leverage 3D ground-truth (GT) data for semi-supervised training. HRN utilizes the entire 3D ground-truth scan to attain both deformation and displacement maps. Ours only takes a sparse set of 3D landmarks.
TABLE I: Comparison to SOTAs.
@@
Deep3DFace [10] Pixel-Face 1.6641
MGCNet [11] Pixel-Face 1.8877
@@
Deep3DFace [10] Bosphorus 1.4777
MGCNet [11] Bosphorus 1.4418
@@
HRN [24] Pixel-Face 1.4061
Ours Pixel-Face 1.4040
Given the scarcity of open sources for this task, we put our full codes (training & test) in supplementary; the pre-trained weight and code will also be provided in our GitHub.
Comparsion on Pixel-Face. The quantitative results of the comparison are shown in Tab. I. Since Deep3DFace [10] and MGCNet [11] did not use 3D landmarks, to be fair, we also provide the results of our model without using 3D landmarks for comparison. As observed, our model (without 3D landmarks) shows relative $5.2\%$ improvement compared to the existing methods. While HRN [24] adopts the complete 3D ground-truth scan to attain the deformation map and displacement map, we only take very light annotation, 101 3D landmarks to constraint the face reconstruction. Ours shows very competitive results with fewer 3D annotation, compared to HRN. The visual comparison is shown in Fig. 5 given 3-view faces. It is evident that our predicted model is more accurate, especially in terms of facial depth estimation in the facial features.
@@
Comparsion on Bosphorus. Since the Pixel-Face dataset is collected based on Asian faces, the samples it contains are limited by region. Moreover, the Pixel-Face dataset is only made under one lighting environment and three fixed camera angles. Taking into account the diversity, we adopt the richer Multi-PIE dataset for training and the Bosphorus dataset for testing. Both datasets contain faces across continents and are collected in various lighting environments and shooting angles. Because Multi-PIE does not provide 3D landmarks, we remove $L_{l\_3d}$ to train DF-MVR for this part. The traditional delaunay algorithm is utilized to generate mesh files from point clouds for the Bosphorus dataset. We set the left and right inputs to 45 degrees for comparison. More angle robustness experiments can be later seen in Tab. III. The quantitative and qualitative comparisons are shown in Tab. I and Fig. 6. Our model (without 3D landmarks) exhibits relative $3.0\%$ improvement compared to the existing methods.
@@
V-C Ablation Study
In order to verify the effectiveness of the proposed feature fusion backbone and the mask mechanism we designed, we perform more ablation experiments, as shown in Tab. II. First, from v1, v2, and v6, it can be found that the multi-view feature fusion network we designed is superior to traditional CNN and Unet in this task. Then, the results of v2 and v6 hint that the multi-layer feature interaction in the feature extraction stage is better than the direct concatenation of features at the end. Through the RMSE of v3 and v6, it is clear that RedNet performs better than ResNet in this task. For v4, we not only remove the mask loss but also the face mask $\mathbf{I}_{A}$, $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$, which are concatenated to the original image. By comparing v5 and v6, we can observe that the face mask mechanism promotes the network to generate a higher-precision model. Moreover, we verify the face mask and mask loss separately. Mask loss has a much smaller effect on the accuracy of the model than face mask. While the qualitative result of mask loss is positive for the adjustment of the model during training, we retain this part for reference. Finally, we remove $L_{lan\_3d}$, which means that our model can be trained with only three multi-view images without any 3D label (as denoted as v5). The result also shows that our model is accurate and stable.
TABLE II: Ablation Study
@@ V-C Ablation Study
v1 CNN RedNet50 Y Y 1.5415
v2 Unet RedNet50 Y Y 1.5207
v3 Ours ResNet50 Y Y 1.5206
v4 Ours RedNet50 N Y 1.4656
v5 Ours RedNet50 Y N 1.5770
v6 Ours RedNet50 Y Y 1.4040
V-D Robustness Analysis
Figure 7: Robustness analysis.
Multi-PIE and Bosphorus are collected by different cameras. DF-MVR uses Multi-PIE for training and achieves good test results on Bosphorus, which implies that our model has a certain robustness to camera device changes. Moreover, we further explore the robustness of DF-MVR for shooting angles. We fix the $\mathbf{I}_{B}$ and $\mathbf{I}_{C}$ face angles while adjusting the $\mathbf{I}_{A}$ face angle. Tab. III and Fig. 7 show that our model consistently achieves the best results in different shooting angles, implying that our model exhibits the best robustness. Random means we choose the degree from $10^{\circ}$, $20^{\circ}$, $30^{\circ}$ and $45^{\circ}$ randomly. The light environment is also random in this experiment.
TABLE III: Robustness Analysis on Bosphorus dataset.
@@ V-D Robustness Analysis
Deep3DFace [10] 1.5030 1.4975 1.4924 1.4777 1.4905
MGCNet  [11] 1.4419 1.4412 1.4413 1.4418 1.4416
@@
VI Conclusion
In this paper, we design a novel end-to-end weakly supervised multi-view 3D face reconstruction network that exploits multi-view encoding to a single decoding framework with skip connections, able to extract, integrate, and compensate deep features with attention. In addition, we develop a multi-view face parse network to learn, identify, and emphasize the critical common face area. Combining pixel-wise photometric loss, mask loss, and landmark loss, we complete the weakly supervised training. Extensive experiments verify the effectiveness of our model. In future work, we plan to extend our framework by incorporating NeRF and Gaussian Splitting to enhance the representation capabilities of our method.